OpenStudy (anonymous):

What are the approximate solutions of 5x2 − 7x = 1 rounded to the nearest hundredth? No real solutions x ≈ −0.65 and x ≈ 7.65 x ≈ −1.53 and x ≈ 0.13 x ≈ −0.13 and x ≈ 1.53

OpenStudy (anonymous):

ok you should solve this using Quadratic equation. \[\huge 5x^2-7x-1=0\] here a=5,b=-7,c=-1

OpenStudy (anonymous):

@rebecca1233 did you get it?

OpenStudy (anonymous):

OK, quick and dirty, without a huge amount of calculation. The quadratic function opens upwards (leading coefficient positive) and the y-intercept is negative. That means there will be two distinct real roots, so the first answer out. The axis of symmetry for the graph is x=-b/2a=7/10, so the solutions have to be equal distance above and below that x value. The only one of your given answers that meets that specification is the last one.

OpenStudy (anonymous):

Think about the way the graph looks, and you can narrow the field of possible solutions quickly. Of course, you can grind it out as shown very nicely by sami-21.